The adjacent sides of a rectangle differ by 5 feet.If the area of the rectangle is 84 square feet, what is the lenght of the smaller side?
Answer by mananth(16946)

let the smaller side be x
the larger side will be x+5
Area = 84 ft^2
x(x+5)=84
x^2+5x=84
x^2+5x-84 =0
x^2+12x-7x-84=0
x(x+12)-7(x+12)=0
(x+12)(x-7)=0
x = 7 the smaller length
the larger length will be 12
...
m.ananth@hotmail.ca